Fermi National Laboratory
 

Operation Summary for November 4, 2011 - November 7, 2011

 

NuMI reported receiving 1.04E19 protons on target for the period from 10/31/11 to 11/7/11. This set a record for being the highest number of protons on target for a week.

Friday November 4

The day shift began with Operations monitoring the Main Injector (MI), Linac, and Booster, with Switchyard (SY) personnel conducting a study, and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, at 8:08 AM, Neutron Therapy Facility (NTF) personnel began preparing to treat two patients later in the morning. At 10:55 AM, the SY beam permit tripped off due to beam problems during a study; operators and others investigated the problem. The NuMI beam permit tripped off when a quadrupole power supply (QF114) tripped off at 1:21 PM; a few minutes later NuMI went down again due to a BPM not polling. Operators rebooted the BPM (NBPM65) system and resumed sending beam to NuMI at 1:40 PM. The SY study ended at 2:54 PM. Booster personnel also conducted a study.

During the evening shift, the Booster personnel ended their study at 8:10 PM. Operators reported MI beam position problems and adjusted the tunes.

Saturday November 5

The midnight shift began with Operations sending beam NuMI, and MiniBooNE.

During the midnight shift, at 12:20 AM, NuMI went down again due to NBPM65 not polling; operators rebooted the system. Operators tuned MI at 2:21 AM.

During the day shift, operators tuned MI at 9:29 AM.

During the evening shift, operators tuned the 750 KeV beam line at 7:41 PM. At 4:29 PM, a Linac RF station (LRF2) tripped off for the first of two times in about five hours; it reset.

Sunday November 6

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 12:57 AM, LRF1 tripped off for the first of two times in about an hour; it reset. At 1:51 AM, LRF2 tripped off for the first of three times in about two hours; it reset. At 6:05 AM, operators reported a liquid nitrogen alarm; operators conferred with a Mechanical person and then searched for a leak. Operators found the leak, fixed it, and at the Mechanical person’s request, contacted the “Air Products” company and ordered an immediate refill for the A0 liquid nitrogen Dewars.

During the day shift, at 9:01 AM, the I- Source ion pump #4 tripped off; it reset. At 9:44 AM, the I- Source ion pumps 1, 2, and 3 tripped off; they all reset. An “Air Products” nitrogen truck filled the A0 Dewars at 9:49 AM. Operators conducted a tune of the 400 MeV line at 11:08 AM.

During the evening shift, Operations monitored the Accelerator Complex.

Monday November 7

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 1:52 AM, a Linac Klystron RF station (KRF5) tripped for the first of three times in one minute; it reset. At 7:43 AM, an MI abort gap kicker tripped off and would not reset; the Operations MI Specialist instructed operators to run without any of the gap cleaning kickers until a Kicker person could investigate. Operators reported minor downtimes due to Linac and Booster RF trips.

Operations Coordinator (OpsCo)

The OpsCo approved of the plans by Booster and MI personnel to conduct studies. NuMI reported that on Sunday they set a new record for the highest number of protons on target for 24 hours of 1.56E18.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 11/07/2011    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory